OpenStudy (anonymous):

write the first five terms of the sequence. find the common difference of tge sequence. write the nth term of the sequence as a function of n. a1=200 a = a -10 k+1 k

OpenStudy (amistre64):

how far have you gotten?

OpenStudy (amistre64):

and, it simpler to define this as ak = a{k-1} - 10, that wayt he kth term doesnt need to be mentally altered

OpenStudy (cwrw238):

if i've understood it correctly 2nd term = a2 = a(2-1) - 10 = 200-10 = 190 CD = -10 first 5 terms are 200,190,180,170,160

OpenStudy (cwrw238):

nth term = 200 - 10n

OpenStudy (amistre64):

thats the issue with starting a seq at n=1, it messes with the count

OpenStudy (amistre64):

a1 = 200; yet, 200-10(1) = 190